# SVTRv2: CTC Beats Encoder-Decoder Models in Scene Text Recognition

### v1

@@ 7 SVTRv2 Variants
8 More detail of real-world datasets
For English recognition, we train models on real-world datasets, from which the models exhibit stronger recognition capability [4, 25, 37]. There are three large-scale real-world training sets, i.e., the Real dataset [4], REBU-Syn [37], and the training set of Union14M-L (U14M-Train) [25]. However, as shown in Fig. 6 and Tab. 8, the former two significantly overlap with Union14M-Benchmarks, thus not suitable for model training. Surprisingly, U14M-Train is also overlapped with Union14M-Benchmarks, in nearly 6.5k text instances across the seven subsets. It means the models trained based on U14M-Train suffer from data leakage when tested on Union14M-Benchmarks,, thus the results reported by [25] should be updated. To this end, we create a filtered version of Union14M-L training set, termed as U14M-Filter, by filtering these overlapping instances. This new dataset is used to train SVTRv2 and other 24 methods we reproduced.
9 More detail of Inference Time
In term of the inference time, we do not utilize any acceleration framework and instead employ PyTorch’s dynamic graph mode on one NVIDIA 1080Ti GPU. We first measure the inference time for 3,000 images with a batch size of 1, calculating the average inference time for each text length. We then computed the arithmetic mean of the average time across all text lengths to determine the overall inference time of the model. Algorithm 1 details the process of measuring inference time.
10 Results when trained on synthetic datasets
Previous research typically follows a typical evaluation protocol, where models are trained on synthetic datasets and validated using six widely recognized real-world benchmarks. Building on this protocol, we trained the SVTRv2 model on synthetic datasets. In addition to evaluating SVTRv2 on the six common benchmarks, we also assess its performance on challenging benchmarks, i.e. Union14M-Benchmark, offering a comprehensive understanding of the model’s generalization capabilities. For methods that have not reported performance on challenging benchmarks, we conducted additional evaluations using their publicly available models and present these results for comparative analysis. As illustrated in Tab. 9, models trained on synthetic datasets exhibit notably reduced performance compared to those trained on large-scale real-world datasets (see Tab. 3). This performance drop is particularly pronounced on challenging benchmarks. These findings emphasize the critical importance of real datasets in improving recognition accuracy for challenging text scenarios.
Despite the challenges associated with synthetic datasets, SVTRv2 exhibits superior performance across both average accuracy metrics, surpassing the previously best-performing method by 0.22% and 2.73%, respectively. On irregular text benchmarks, such as Curve and Multi-Oriented, SVTR achieves strong results, largely due to its integrated correction module, which is particularly adept at handling irregular text patterns, even when trained on synthetic datasets. Notably, SVTRv2 achieves a substantial 4.8% improvement over SVTR on the Curve benchmark, further demonstrating its enhanced capacity to address irregular text. Overall, these results demonstrate that, even when trained solely on synthetic datasets, SVTRv2 exhibits strong generalization capabilities, effectively handling complex and challenging text recognition scenarios.
@@ 11 Qualitative Analysis of Recognition Results
The SVTRv2 model achieved an average accuracy of 96% on a standard dataset. To investigate the underlying causes of the remaining 4% of recognition errors, we conducted a detailed analysis of misclassified samples across six standard datasets, as illustrated in Fig. 7 and Fig. 8. While previous research has typically categorized common benchmarks into regular and irregular text. However, these error samples indicates that the majority of incorrectly recognized text is not irregular. This suggests that, under the current training paradigm using large-scale real-world datasets, a more rigorous manual screening process for common benchmarks is warranted.
@@
| IIIT5k | 0 | 16 | 1 | 4 | 21 | 4 |
| SVT | 4 | 4 | 4 | 0 | 12 | 0 |
| ICDAR 2013 | 2 | 2 | 4 | 2 | 10 | 2 |
| ICDAR 2015 | 48 | 19 | 42 | 13 | 122 | 35 |
| SVTP | 7 | 6 | 12 | 7 | 32 | 4 |
| CUTE80 | 0 | 1 | 0 | 0 | 1 | 1 |
@@
Consequently, we identified five primary causes of recognition errors in these samples: (1) blurred, (2) artistic, (3) incomplete text, (4) image text labeling inconsistency, and (5) others. Specifically, the blurring text category encompasses issues such as low resolution, motion blur, or extreme lighting conditions. The artistic text category includes unconventional print fonts, predominantly found in business signage, as well as a proportion of handwritten text. Incomplete text refers to characters obscured by other objects or missing due to improper cropping, where missing information must be inferred from context. Image text labeling inconsistency means that there is an error in the given text label or there are some characters with phonetic symbols. As shown in Tab. 10, after excluding samples where the errors were due to labeling inconsistency, the remaining errors could be attributed to blurred (30.81%), artistic (24.24%), and incomplete text (31.82%), respectively. This classification allows us to conclude that SVTRv2’s recognition performance, particularly in complex scenarios involving blurred, artistic, or incomplete text, requires further enhancement.
@@ 12 Standardized Model Training Settings
The optimal hyperparameters for training different models often vary and are not universally fixed. However, critical factors such as training epochs, data augmentation techniques, input size, data type, and evaluation protocols have a substantial influence on model accuracy. To ensure fair and unbiased performance comparisons across models, these factors must be strictly standardized. Accordingly, we adopt a uniform training and evaluation setting, as shown in Table 11, to maintain consistency across all settings while simultaneously enabling each model to achieve its best possible accuracy.
@@
| Data Type | All models are trained with mixed precision. |

### v2

@@ 10 Results when Trained on Synthetic Datasets
| CPPD [13] | TPAMI 2025 | SVTR-B | 97.6 | 95.5 | 98.2 | 87.9 | 90.9 | 92.7 | 93.80 | 65.5 | 18.6 | 56.0 | 61.9 | 71.0 | 57.5 | 65.8 | 56.63 | 26.8 | ||||||||||||||
| IGTR-AR [14] | TPAMI 2025 | SVTR-B | 98.2 | 95.7 | 98.6 | 88.4 | 92.4 | 95.5 | 94.78 | 78.4 | 31.9 | 61.3 | 66.5 | 80.2 | 69.3 | 67.9 | 65.07 | 24.1 | ||||||||||||||
| SMTR [12] | AAAI 2025 | FocalSVTR | 97.4 | 94.9 | 97.4 | 88.4 | 89.9 | 96.2 | 94.02 | 74.2 | 30.6 | 58.5 | 67.6 | 79.6 | 75.1 | 67.9 | 64.79 | 15.8 | ||||||||||||||
@@ 11 Qualitative Analysis of Recognition Results
The SVTRv2 model achieved an average accuracy of 96.57% on Com (see Tab. 3). To investigate the underlying causes of the remaining 3.43% of recognition errors, we conducted a detailed analysis of the misclassified samples, as illustrated in Fig. 7 and Fig. 8. While previous research has typically categorized Com into regular and irregular text. However, these error samples indicate that the majority of incorrectly recognized text is not irregular. This suggests that, under the current training paradigm using large-scale real-world datasets, a more rigorous manual screening process is warranted for common benchmarks.
@@
| IIIT5k [33] | 0 | 16 | 1 | 4 | 21 | 4 |
| SVT [44] | 4 | 4 | 4 | 0 | 12 | 0 |
| ICDAR 2013 [27] | 2 | 2 | 4 | 2 | 10 | 2 |
| ICDAR 2015 [26] | 48 | 19 | 42 | 13 | 122 | 35 |
| SVTP [35] | 7 | 6 | 12 | 7 | 32 | 4 |
| CUTE80 [1] | 0 | 1 | 0 | 0 | 1 | 1 |
@@
Based on this one-by-one manual viewing, we identified five primary causes of recognition errors: (1) blurred, (2) artistic, (3) incomplete text, (4) others, and (5) image text labeling errors (Labelerr). Specifically, the blurring text includes issues such as low resolution, motion blur, or extreme lighting conditions. The artistic text category refers to unconventional fonts, commonly found in business signage, as well as some handwritten text. Incomplete text arises when characters are obscured by objects or lost due to improper cropping, requiring contextual inference. Image text labeling errors occur when the given text labels contain inaccuracies or include characters with phonetic symbols. As shown in Tab. 11, after excluding samples affected by labeling inconsistencies, the remaining recognition errors primarily stemmed from blurred (30.81%), artistic (24.24%), and incomplete text (31.82%). This result highlights that SVTRv2’s recognition performance needs further improvement, particularly in handling complex scenarios involving these challenging text types.
@@ 12 Standardized Model Training Settings
The optimal hyperparameters for training different models vary and are not universally fixed. However, key factors such as training epochs, data augmentations, input size, and evaluation protocols significantly influence model accuracy. To ensure fair and unbiased performance comparisons, we standardize these factors across all models, as outlined in Tab. 12. This uniform training and evaluation framework ensures consistency while allowing each model to approach its best accuracy. To maximize fairness, we conducted extensive hyperparameter tuning for model-specific settings, including the optimizer, learning rate, and regularization strategies. This rigorous optimization led to significant accuracy improvements of 5–10% for most models compared to their default configurations. For instance, MAERec’s accuracy increased from 78.6% to 85.2%, demonstrating the effectiveness of training settings. These improvements underscore the reliability of our results and highlight the importance of carefully optimizing hyperparameters for meaningful model comparisons.